(D^2+1)y=x^3+e^xcosx
1
Expert's answer
2021-07-19T05:48:46-0400
"y''+y=x^3+e^x\cos x"

Homogeneous equation


"y''+y=0"

Characteristic equation


"r^2+1=0"

"r=\pm i"

The general solution of the homogeneous equation is


"y_h=c_1\cos x+c_2\sin x"

Find the partial solution of the nonhomogeneous equation in the form


"y_p=Ax^3+Bx^2+Cx+D+Ee^x\cos x+Fe^x\sin x"

"y_p'=3Ax^2+2Bx+C+Ee^x\cos x-Ee^x\sin x"

"+Fe^x\sin x+Fe^x\cos x"


"y_p''=6Ax+2B+Ee^x\cos x-Ee^x\sin x"

"-Ee^x\sin x-Ee^x\cos x+Fe^x\sin x+Fe^x\cos x"

"+Fe^x\cos x-Fe^x\sin x"

Substitute


"6Ax+2B-2Ee^x\sin x+2Fe^x\cos x"

"+Ax^3+Bx^2+Cx+D+Ee^x\cos x+Fe^x\sin x"

"=x^3+e^x\cos x"

"x^3: A=1"

"x^2: B=0"

"x^1: 6A+C=0"

"x^0: 2B+D=0"

"e^x\cos x:E+2F=1"

"e^x\sin x:-2E+F=0"

"A=1, B=0, C=-6, D=0, E=\dfrac{1}{5}, F=\dfrac{2}{5}"

"y_p=x^3-6xe^x+\dfrac{1}{5}\cos x+\dfrac{2}{5}e^x\sin x"

The general solution of the given nonhomogeneous equation is


"y=c_1\cos x+c_2\sin x"




"+x^3-6xe^x+\dfrac{1}{5}\cos x+\dfrac{2}{5}e^x\sin x"